1. The value of KC for the thermal decomposition of hydrogen sulfide, shown below, is 2.2 × 10-4 at 1400 K.
2H2S(g) ⇌ 2H2(g)+ S2(g)
A sample of gas in which [H2S] = 3.60 M is heated to 1400 K in a sealed vessel. After chemical equilibrium has been achieved, what is the value of [H2S]? Assume no H2 or S2 was present in the original sample. answer in M
1
Expert's answer
2020-04-06T11:48:22-0400

"2H_2S(g) \leftrightarrow 2H_2(g) + S_2(g)"

Make ICE table

"[H_2S]" "[H_2]" "[S_2]"

Initial 3.60 0 0

Change -x +x +0.5x

Equilibrium 3.60-x x 0.5 x


"K_c = \frac{[H_2]^2[S_2]}{[H_2S]^2}"


"2.2\times 10^{-4} = \frac{(x^2)(0.5x)}{(3.60-x)^2}"


make an assumption that 3.60-x = 3.60


"2.2\times10^{-4} = \frac{(x^2)(0.5x)}{(3.60)^2}"


"x=0.179"


"[H_2S] = 3.6-0.179 = 3.421 M"